OpenStudy (anonymous):

Solve the following equation. Please show all of your work. 7+ 5/7 x =-13

OpenStudy (gw2011):

7 + (5/7)x = -13 (5/7)x = -13-7 (5/7)x = -20 x = (-20)(7/5) x = -28